A compound forms hexagonal close-packed structure. What is the total number of voids in 0.5 mol of it? How many of these are tetrahedral voids?

Grade:

5 Answers

Gaurav
9 years ago
Number of close-packed particles = 0.5×6.022×1023= 3.011×1023

Therefore, number of octahedral voids = 3.011×1023

And, number of tetrahedral voids = 2×3.011×1023= 6.022×1023

Therefore, total number of voids = 3.011×1023+ 6.022×1023= 9.033×1023
sanjib malakar
5 years ago
Number of atoms in close packaging = 0.5 mol
1 has 6.022 ×10 23  particles
So that
Number of close−packed particles
= 0.5 × 6.022 × 1023 = 3.011 × 1023
Number of tetrahedral voids = 2 × number of atoms in close packaging
Plug the values we get
Number of tetrahedral voids = 2 × 3.011 × 1023
                                                  = 6.022 × 1023
 Number of octahedral voids = number of atoms in close packaging
So that
Number of octahedral voids = 3.011 × 1023
Total number of voids = Tetrahedral void + octahedral void
                                       = 6.022 × 1023 + 3.011 × 1023 
                                       = 9.03 × 1023
Vyshnav
5 years ago
Number of atoms in close packaging = 0.5 mol
1 has 6.022 ×1023particles
So that
Number of close−packed particles
= 0.5 × 6.022 × 1023 = 3.011 × 1023
Number of tetrahedral voids = 2 × number of atoms in close packaging
Plug the values we get
Number of tetrahedral voids = 2 × 3.011 × 1023
                                                  = 6.022 × 1023
 Number of octahedral voids = number of atoms in close packaging
So that
Number of octahedral voids = 3.011 × 1023
Total number of voids = Tetrahedral void + octahedral void
                                       = 6.022 × 1023 + 3.011 × 1023 
                                       = 9.03 × 1023
Life Is A PARADOX
5 years ago
 
We clearly know that  the total no. of particles
( let n),  n = no. of moles  * avagadro number
                 = 0.5 * 6.022 * 10^23
                 = 3.011 * 10^23 particles 
 
Also we know that that octahedral voids are equal 
To the number of particles of the packed structure while the tetrahedral voids are generally the double of the number of particles of the packed structure. 
 
Hence 
Octahedral voids =  n = 3.011 * 10^23 voids (O) 
And 
Tetrahedral voids = 2n  = 2  * 3.011 * 10^23
                                     = 6.022 * 10^23 voids (T) 
 
Total no.  of voids = (O) + (T) 
                                = 6.022 * 10^23 + 3.011 * 10^23
                                = 9.033 * 10^ 23 voids 
Nisha shaji
2 years ago
A compound forms hcp structure. What is the total number of voids in 0.5 mole of it? How many of these are tetrahedral voids?